# NeLLCom-X: A Comprehensive Neural-Agent Framework to Simulate Language Learning and Group Communication

### v1

@@ 5 Replicating the Trade-off with Full-fledged Agents
After SL, our agents have successfully learnt both languages but no regularization happens, as expected. By contrast, the results of self-play averaged over each 50-agent set indicate that both languages progressively lose markers. Crucially, the fixed-order language does so faster than the flexible one, where markers are often necessary for agent/patient disambiguation. In sum, self-play in NeLLCom-X results in very similar trends as the simple NeLLCom setup, confirming the emergence of a human-like order/marking trade-off Fedzechkina et al. (2017). Detailed replication results are provided in Appendix B. Here, we report communication success during self-play and production preferences at the end of self-play for the flexible language (Table 2, left column). Self-understanding increases through RL leading to a much more informative language, while production preferences reveal that this spans from an overall decrease in order entropy with marking proportion remaining almost the same on average (solid circle). While some agents approach the optimal points of fixed-order/no-marking (bottom-left corner) or flexible-order/full-marking (top-right), the large variability in production preferences suggests many agents settle on less optimized, redundant languages, as also found by Lian et al. (2023).
50s+67m 50s+50m
Comm. success
per turn
@@
Table 2: Two populations of 50 agents engaging in self-play (no interaction) after having learned two flexible-order, optional-marker languages: one with 67% the other with 50% marking. Top row: Average communication success across self-play turns. Bottom row: Production preferences: solid diamonds mark the initial language; each empty circle denotes a full-fledged agent at the end of self-play; solid circles are the average of all agents, with error bars showing standard deviation.
@@
We reconsider here a language design choice of Lian et al. (2023) who, in turn, inherited it from the human study of Fedzechkina et al. (2017). It was recently found that human learners exposed to a fixed-order language with 75% marking tend to regularize by increasing marker use even though this would make the language less efficient Tal et al. (2022). Similarly, the dominant proportion (67%) of marking utterances in our initial languages may push the agents to prefer marking even when it may be a redundant strategy. Hence, we propose that a more balanced distribution of 50% markers and 50/50% word order may be a better choice to reveal the intrinsic preferences of the learners, if there are any, without biasing them to regularize markers. Results in Table 2 (right column) show that this language has overall lower communicative success, as expected given the higher amount of ambiguous sentences. However, success increases substantially during interaction while production preferences reveal a larger variability in solutions including those with more fixed order and less markers. We use this more neutral combination as the default language in all remaining experiments.
@@ 6.1 Speakers of Different Languages
$\alpha_{base}\leftrightarrow\alpha_{base}$ $\alpha_{base}\leftrightarrow$ 80s+20m $\alpha_{base}\leftrightarrow$ 20s+20m $\alpha_{base}\leftrightarrow$ 50s+80m $\alpha_{base}\leftrightarrow$ 80s+50m
Comm. success
per turn
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Marker use by
order (SOV/all)
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 3: Interactive communication between different language speakers. The first agent is always trained on 50s+50m ($\alpha_{base}$). Each experiment is repeated with 50 agent pairs.
@@
As explained in Section 3.3, each agent performs a turn of self-play after completing $\sigma=10$ turns of interactive communication, based on preliminary experiments. We compare this to a setup where no self-play is performed during interaction ($\sigma=\inf$), in the case where two agents start from a state of poor mutual understanding due to limited marking and strongly diverging order preferences (80s+20m vs. 20s+20m). As shown in Table 4, disabling self-play leads to extremely low self-understanding even though communication between the two agents is successful. To explain this result, we inspect the production preferences of individual agent pairs and find that many regularize their language in opposite directions (e.g. dominant SOV vs. dominant OSV, both with no markers), indicating a total decoupling of the speaking and listening ability. Thus, we confirm that embedding tying alone does not allow for a realistic interaction simulation, making self-play necessary in our framework.
@@
Table 4: Impact of self-play during interaction in pairs of agents speaking 80s+20m and 20s+20m respectively. Each experiment is repeated with 20 agent pairs.
@@ 6.2 Effect of Group Size
Here we move back to a setup where all agents are trained on the same neutral and unstable initial language (50s+50m), but this time they interact in groups of different sizes (2, 4, 8, 20) using the standard self-play frequency ($\sigma=10$). To make results comparable, we ensure the total number of interactive turns per agent is the same ($\approx$200) in all setups, by setting $comm\_round$ to 100, 34, 15, and 6 respectively. A total of 200 agents are trained in each group size setting.444100 runs of group of 2, 50 of 4, 25 of 8, and 10 of 20. See all group-specific training details in Appendix C.
groups of 2 groups of 4 groups of 8 groups of 20
Comm. success
per turn
Marker use by
order entropy
@@
Table 5: Interactive communication in groups of same-language speakers (50s+50m). Bottom row: Group-level production preferences (each point is a group) and Spearman’s correlation $\rho$ between marker use and order entropy.
@@
Table 5 (top row) shows similar learning curves for all group sizes, demonstrating that communication is successful even in larger groups. In all cases, interactive and self-communication test accuracy start low (25%), but agents collaborate and end up between 60% and 80% success at $\mathsf{inter\_turn}$ = 100.
@@
For production preferences, we plot proportion of marking by order entropy as we are again interested in order flexibility rather than the specific order chosen by the agents (Table 5, bottom row). Here, each circle denotes the average production preferences of an entire group, as opposed to those of a single agent. When comparing results across different group sizes, we see that the variability observed in self-playing agents (Section 5) including less optimal and redundant strategies, gets smaller as group size increases. The average entropy in groups of 8 and 20 is also lower than in groups of 4 or 2. In the group setting, an agent’s choice to use a marker does not only depend on its own order entropy but on that of the entire group. As a measure of the order/marking trade-off at group level, we therefore calculate Spearman’s correlation ($\rho$) between order entropy and marker use, both computed over all (categorizable) utterances produced by all agents in a group. As shown in Table 5, $\rho$ steadily increases with group size from relatively weak (0.32) in pairs to strong (0.73) in groups of 20. This confirms that pairs, like self-playing agents, still often settle on redundant strategies, while larger groups develop more optimized languages in which stronger order consistency at the group level leads to a drop in marker use, confirming the emergence of the trade-off also at the group level. [^5] Even when trained for much longer, the results of pairs remain similar, suggesting they indeed settle on less optimized solutions which is not overcome simply by more interactions (e.g. 200 rounds, $\rho$ = 0.33). See Appendix D.
@@ 7 Discussion and conclusion
We introduced NeLLCom-X, a framework for simulating neural agent language learning and communication in groups, starting from pre-defined languages. Agents in this framework display the cognitively plausible property of interchangeability Hockett (1960), by which anything they can understand, they can say and vice versa, while also having the ability to align to other individuals. We replicated an earlier finding by Lian et al. (2023) and showed that a word-order/case-marking trade-off still appears with the adjusted full-fledged agent architecture. Subsequently, we simulated interactions between agents trained on different languages. We found that pairs quickly adapt their utterances towards a mutually understandable language and that the neutral language drifts in different directions depending on the preferences of the other agent. Moreover, agents converge on a shared language faster, and reach higher accuracy in cases where one of the two agents has a stronger word order preference. We then assessed the effect of performing self-play during interactive communication and found it necessary to ensure our full-fledged agents continue to understand themselves, while also realistically adapting to other individuals. Lastly, we studied group dynamics and found that NeLLCom-X agents manage to establish a successful communication system even in larger groups (up to size 20). Moreover, we generally see a larger entropy reduction in the languages developed by larger groups as compared to the languages used by pairs of agents. This finding aligns with previous work on group-level emergent communication, where it was shown that groups developed less idiosyncratic languages than pairs Tieleman et al. (2019) as well as with human experiments which demonstrated more systematic languages to emerge in larger groups Raviv et al. (2019). In our simulations, pairs and smaller groups sometimes settle on less optimized and partly still redundant solutions, while large groups end up with more efficient communication systems. We discuss the current limitations of our work in Supplementary Material.
@@
Although the use of miniature artificial languages in our work allows for easily interpretable results due to abstractions and simplifications that are hard to achieve with natural human languages, the languages used currently are very small. This may limit the possibility of drawing conclusions beyond proof-of-concept demonstrations. Future work should increase the size and complexity of the languages to see if results hold on a larger scale.
@@
Interaction between individuals in groups is not the only population factor that shapes language. Linguistic structure is shaped by both interaction and learning Kirby et al. (2015). Especially when languages are learned and transmitted to subsequent generations repeatedly, even small inductive biases may have a large effect on emerging properties Thompson et al. (2016). We plan to augment NeLLCom-X with iterated learning so that new agents learn from the utterances of others and become teachers to agents in the next generation.
@@ Appendix A More Details about NeLLCom
(a) 100s+67m (b) 100s+50m (c) 50s+67m (d) 50s+50m
(1) Comm. success
@@
Table 6: Replicating the results from Lian et al. (2023) with NeLLCom-X full-fledged self-communicating agents with fixed-order (a) and flexible order (c) languages. Comparing the original results with a new, more neutral, initial languages with 50% markers in (b) and (d).
@@ B.2 Results
Starting from the initial marker proportion (66.7%), fixed-order language learners start to drop the marker (50% at round 60) during self-communication while maintaining high understandability (95%) (Table 6 (1a) and (4a)). This aligns with the results of Lian et al. (2023).
@@
The self-communication accuracy in the flexible-order language (Table 6 (1c)) starts from a relatively low success rate as expected, but increases with more communication rounds. In particular, agents exceed the communication success they had achieved at the end of SL on new meanings and finally reach a much higher accuracy on new meanings at the end of self-communication (around 75%) comparing to the communication success they had achieved at the end of SL.
The average ordering and marking proportions also show that flexible-order language self-communication results in a very similar pattern as was found by Lian et al. (2023): (i) The average word order production (Table 6 (2c)) shows a strong preference for OSV, (ii) Although the overall marking system ends with a similar marker proportion as the initial condition (Table 6 (4c)), i.e., the proportion of with-marker utterances is twice the proportion of no-marker utterances, we can see a clear shift to conditional marking (Table 6 (3c)) with an asymmetric use of markers: at round 60, the marker proportion on utterances with OSV order (70%) remains similar to the initial proportion (66.7%), while the proportion of markers use with SOV drops to 35%. This order preference and asymmetric marking system align with the flexible-order language results of Lian et al. (2023).
@@
50% marking in initial language
As described in Section 5, the initial proportion of marker use of 67%, which was used in Lian et al. (2023) and inherited from Fedzechkina et al. (2017), may create a bias for the agents to regularize towards more marker use, settling on more redundant languages. We therefore switched to the more neutral value of 50% markers in the initial language. In the first four rows of Table 6, the self-communication results of this new setting can be directly compared to the original set-up. As expected, markers are dropped more rapidly in the fixed-order 50% marker language than in the 67% marker language (Table 6 (3a) versus Table 6 (3b)). In the flexible-order languages, agents trained on the 67% marker language mostly kept using the marker, even though they also developed a clear preference for one word order, resulting in redundant strategies. With 50% markers in the initial language, however, agents drop the marker when they develop a word order preference despite being trained on a flexible word order language (Table 6 (3c) versus Table 6 (3d)).
@@ Appendix D Additional Group Experimennts
Table 8 shows the effect of longer training on the production preferences of pairs of same-language speakers (50s+50m). Production preferences (bottom row) do not change much after 100 additional turns (right column), and the correlation $\rho$ increases only marginally from 0.32 to 0.33.
groups of 2 (100 rounds) groups of 2 (200 rounds)
Comm. success
per turn
[Uncaptioned image] [Uncaptioned image]
Marker use by
order entropy
[Uncaptioned image] [Uncaptioned image]
Table 8: Interactive communication in pairs of same-language speakers (50s+50m): Production preferences (bottom row) do not change much when training for 200 rounds (right column) instead of 100 (left).

### v2

@@ 5 Replicating the Trade-off with Full-fledged Agents
After SL, our agents have successfully learnt both languages but no regularization happens, as expected. By contrast, the results of self-play averaged over each 50-agent set indicate that both languages progressively lose markers. Crucially, the fixed-order language does so faster than the flexible one, where markers are often necessary for agent/patient disambiguation. In sum, self-play in NeLLCom-X results in very similar trends as the simple NeLLCom setup, confirming the emergence of a human-like order/marking trade-off Fedzechkina et al. (2017). Detailed replication results are provided in Appendix B. Here, we report communication success during self-play and production preferences at the end of self-play for the flexible language (Section 5, top row). Self-understanding increases through RL leading to a much more informative language, while production preferences reveal that this spans from an overall decrease in order entropy with marking proportion remaining almost the same on average (solid circle). While some agents approach the optimal points of fixed-order/no-marking (bottom-left corner) or flexible-order/full-marking (top-right), the large variability in production preferences suggests many agents settle on less optimized, redundant languages, as also found by Lian et al. (2023).
Communicative success per turn Marker use by order entropy
50s+67m
@@
Figure 2: Two populations of 50 agents engaging in self-play (no interaction) after having learned two flexible-order, optional-marker languages: one with 67% the other with 50% marking. Left column: Average communication success across self-play turns. Right column: Production preferences: solid diamonds mark the initial language; each empty circle denotes a full-fledged agent at the end of self-play; solid circles are the average of all agents, with error bars showing standard deviation.
@@
We reconsider here a language design choice of Lian et al. (2023) who, in turn, inherited it from the human study of Fedzechkina et al. (2017). It was recently found that human learners exposed to a fixed-order language with 75% marking tend to regularize by increasing marker use even though this would make the language less efficient Tal et al. (2022). Similarly, the dominant proportion (67%) of marking utterances in our initial languages may push the agents to prefer marking even when it may be a redundant strategy. Hence, we propose that a more balanced distribution of 50% markers and 50/50% word order may be a better choice to reveal the intrinsic preferences of the learners, if there are any, without biasing them to regularize markers. Results in Section 5 (bottom row) show that this language has overall lower communicative success, as expected given the higher amount of ambiguous sentences. However, success increases substantially during interaction while production preferences reveal a larger variability in solutions including those with more fixed order and less markers. We use this more neutral combination as the default language in all remaining experiments.
@@ 6.1 Speakers of Different Languages
Communicative success per turn Marker use by order (SOV/all)
$\alpha_{b}\leftrightarrow\alpha_{b}$
[Uncaptioned image] [Uncaptioned image]
$\alpha_{b}\leftrightarrow$ 80s+20m
[Uncaptioned image] [Uncaptioned image]
$\alpha_{b}\leftrightarrow$ 20s+20m
[Uncaptioned image] [Uncaptioned image]
$\alpha_{b}\leftrightarrow$ 50s+80m
[Uncaptioned image] [Uncaptioned image]
$\alpha_{b}\leftrightarrow$ 80s+50m
[Uncaptioned image] [Uncaptioned image]
Figure 3: Interactive communication between different language speakers. The first agent is always trained on 50s+50m ($\alpha_{b}$). Each experiment is repeated with 50 agent pairs.
@@
As explained in Section 3.3, each agent performs a turn of self-play after completing $\sigma=10$ turns of interactive communication, based on preliminary experiments. We compare this to a setup where no self-play is performed during interaction ($\sigma=\inf$), in the case where two agents start from a state of poor mutual understanding due to limited marking and strongly diverging order preferences (80s+20m vs. 20s+20m). As shown in Section 6.1, disabling self-play leads to extremely low self-understanding even though communication between the two agents is successful. To explain this result, we inspect the production preferences of individual agent pairs and find that many regularize their language in opposite directions (e.g. dominant SOV vs. dominant OSV, both with no markers), indicating a total decoupling of the speaking and listening ability. Thus, we confirm that embedding tying alone does not allow for a realistic interaction simulation, making self-play necessary in our framework.
@@
Figure 4: Impact of self-play during interaction in pairs of agents speaking 80s+20m and 20s+20m respectively. Each experiment is repeated with 20 agent pairs, and the average communication per turn is shown.
@@ 6.2 Effect of Group Size
Here we move back to a setup where all agents are trained on the same neutral and unstable initial language (50s+50m), but this time they interact in groups of different sizes (2, 4, 8, 20) using the standard self-play frequency ($\sigma=10$). To make results comparable, we ensure the total number of interactive turns per agent is the same ($\approx$200) in all setups, by setting $comm\_round$ to 100, 34, 15, and 6 respectively. A total of 200 agents are trained in each group size setting.444100 runs of group of 2, 50 of 4, 25 of 8, and 10 of 20. See all group-specific training details in Appendix C. In this paper, we only consider fully connected communication graphs and fix the total amount of trained agents to enable comparison. We leave an exploration of other group communication factors, such as density and connectivity, to future work.
Communicative success per turn Marker use by order entropy
groups of 2
[Uncaptioned image] [Uncaptioned image]
groups of 4
[Uncaptioned image] [Uncaptioned image]
groups of 8
[Uncaptioned image] [Uncaptioned image]
groups of 20
[Uncaptioned image] [Uncaptioned image]
@@
Figure 5: Interactive communication in groups of same-language speakers (50s+50m). Right column: Group-level production preferences (each point is a group) and Spearman’s correlation $\rho$ between marker use and order entropy.
@@
Section 6.2 (left column) shows similar learning curves for all group sizes, demonstrating that communication is successful even in larger groups. In all cases, interactive and self-communication test accuracy start low (25%), but agents collaborate and end up between 60% and 80% success at $\mathsf{inter\_turn}$ = 100.
@@
For production preferences, we plot proportion of marking by order entropy as we are again interested in order flexibility rather than the specific order chosen by the agents (Section 6.2, right column). Here, each circle denotes the average production preferences of an entire group, as opposed to those of a single agent. When comparing results across different group sizes, we see that the variability observed in self-playing agents (Section 5) including less optimal and redundant strategies, gets smaller as group size increases. The average entropy in groups of 8 and 20 is also lower than in groups of 4 or 2. In the group setting, an agent’s choice to use a marker does not only depend on its own order entropy but on that of the entire group. As a measure of the order/marking trade-off at group level, we therefore calculate Spearman’s correlation ($\rho$) between order entropy and marker use, both computed over all (categorizable) utterances produced by all agents in a group. As shown in Section 6.2, $\rho$ steadily increases with group size from relatively weak (0.32) in pairs to strong (0.73) in groups of 20. This confirms that pairs, like self-playing agents, still often settle on redundant strategies, while larger groups develop more optimized languages in which stronger order consistency at the group level leads to a drop in marker use, confirming the emergence of the trade-off also at the group level. [^5] Even when trained for much longer, the results of pairs remain similar, suggesting they indeed settle on less optimized solutions which is not overcome simply by more interactions (e.g. 200 rounds, $\rho$ = 0.33). See Appendix D.
@@ 7 Discussion and conclusion
We introduced NeLLCom-X, a framework for simulating neural agent language learning and communication in groups, starting from pre-defined languages. Agents in this framework display the cognitively plausible property of interchangeability Hockett (1960), by which anything they can understand, they can say and vice versa, while also having the ability to align to other individuals. We replicated an earlier finding by Lian et al. (2023) and showed that a word-order/case-marking trade-off still appears with the adjusted full-fledged agent architecture. Subsequently, we simulated interactions between agents trained on different languages. We found that pairs quickly adapt their utterances towards a mutually understandable language and that the neutral language drifts in different directions depending on the preferences of the other agent. Moreover, agents converge on a shared language faster, and reach higher accuracy in cases where one of the two agents has a stronger word order preference. We then assessed the effect of performing self-play during interactive communication and found it necessary to ensure our full-fledged agents continue to understand themselves, while also realistically adapting to other individuals. Lastly, we studied group dynamics and found that NeLLCom-X agents manage to establish a successful communication system even in larger groups (up to size 20). Moreover, we generally see a larger entropy reduction in the languages developed by larger groups as compared to the languages used by pairs of agents. This finding aligns with previous work on group-level emergent communication, where it was shown that groups developed less idiosyncratic languages than pairs Tieleman et al. (2019) as well as with human experiments which demonstrated more systematic languages to emerge in larger groups Raviv et al. (2019). In our simulations, pairs and smaller groups sometimes settle on less optimized and partly still redundant solutions, while large groups end up with more efficient communication systems.
@@
Although the use of miniature artificial languages in our work allows for easily interpretable results due to abstractions and simplifications that are hard to achieve with natural human languages, the languages used currently are very small. This may limit the possibility of drawing conclusions beyond proof-of-concept demonstrations. Future work should increase the size and complexity of the languages to see if results hold on a larger scale and compare to patterns found in real human languages, such as those reported by Levshina et al. (2023).
@@
Interaction between individuals in groups is not the only population factor that shapes language, but linguistic structure is shaped by both interaction and learning Kirby et al. (2015). Especially when languages are learned and transmitted to subsequent generations repeatedly, even small inductive biases may have a large effect on emerging properties Thompson et al. (2016). We therefore plan to augment NeLLCom-X with iterated learning so that new agents learn from the utterances of others and become teachers to agents in the next generation.
@@ Appendix A More Details about NeLLCom
(1) Comm. success (2) Order use (3) Cond. Marker use (4) Marker user
(a) 100s+67m
@@
Figure 6: Replicating the results from Lian et al. (2023) with NeLLCom-X full-fledged self-communicating agents with fixed-order (a) and flexible order (c) languages. Comparing the original results with a new, more neutral, initial languages with 50% markers in (b) and (d).
@@ B.2 Results
Starting from the initial marker proportion (66.7%), fixed-order language learners start to drop the marker (50% at round 60) during self-communication while maintaining high understandability (95%) (Appendix A (a1) and (a4)). This aligns with the results of Lian et al. (2023).
@@
The self-communication accuracy in the flexible-order language (Appendix A (c1)) starts from a relatively low success rate as expected, but increases with more communication rounds. In particular, agents exceed the communication success they had achieved at the end of SL on new meanings and finally reach a much higher accuracy on new meanings at the end of self-communication (around 75%) comparing to the communication success they had achieved at the end of SL.
The average ordering and marking proportions also show that flexible-order language self-communication results in a very similar pattern as was found by Lian et al. (2023): (i) The average word order production (Appendix A (c2)) shows a strong preference for OSV, (ii) Although the overall marking system ends with a similar marker proportion as the initial condition (Appendix A (c4)), i.e., the proportion of with-marker utterances is twice the proportion of no-marker utterances, we can see a clear shift to conditional marking (Appendix A (c3)) with an asymmetric use of markers: at round 60, the marker proportion on utterances with OSV order (70%) remains similar to the initial proportion (66.7%), while the proportion of markers use with SOV drops to 35%. This order preference and asymmetric marking system align with the flexible-order language results of Lian et al. (2023).
@@ 20 $380=20*(20-1)$ $6=\left\lceil 100/(20-1)\right\rceil$ $10=200/20$
50% marking in initial language
As described in Section 5, the initial proportion of marker use of 67%, which was used in Lian et al. (2023) and inherited from Fedzechkina et al. (2017), may create a bias for the agents to regularize towards more marker use, settling on more redundant languages. We therefore switched to the more neutral value of 50% markers in the initial language. In Appendix A, the self-communication results of this new setting can be directly compared to the original set-up. As expected, markers are dropped more rapidly in the fixed-order 50% marker language than in the 67% marker language (Appendix A (a3) versus Appendix A (b3)). In the flexible-order languages, agents trained on the 67% marker language mostly kept using the marker, even though they also developed a clear preference for one word order, resulting in redundant strategies. With 50% markers in the initial language, however, agents drop the marker when they develop a word order preference despite being trained on a flexible word order language (Appendix A (c3) versus Appendix A (d3)).
@@ Appendix C Training Details for Interactive Communication Experiments
Communicative success per turn Marker use by order entropy
100 rounds
[Uncaptioned image] [Uncaptioned image]
200 rounds
[Uncaptioned image] [Uncaptioned image]
Figure 8: Interactive communication in pairs of same-language speakers (50s+50m): Production preferences (right column) do not change much when training for 200 rounds (bottom row) instead of 100 (top).
@@ Appendix D Additional Group Experimennts
Appendix C shows the effect of longer training on the production preferences of pairs of same-language speakers (50s+50m). Production preferences (right column) do not change much after 100 additional turns (bottom row), and the correlation $\rho$ increases only marginally from 0.32 to 0.33.